OpenStudy (anonymous):

What is an angle between 0 and 2pi that is coterminal with -5pi/6?

jimthompson5910 (jim_thompson5910):

Add 2pi to that given angle to get ???

OpenStudy (anonymous):

210 degrees?

jimthompson5910 (jim_thompson5910):

\[\Large -\frac{5\pi}{6} + 2\pi = ???\] in terms of pi (in radians)

OpenStudy (anonymous):

3.67

jimthompson5910 (jim_thompson5910):

I was aiming more for this \[\Large -\frac{5\pi}{6} + 2\pi = \Large -\frac{5\pi}{6} + 2\pi*\frac{6}{6}\] \[\Large -\frac{5\pi}{6} + 2\pi = \Large -\frac{5\pi}{6} + \frac{12\pi}{6}\] \[\Large -\frac{5\pi}{6} + 2\pi = \Large \frac{-5\pi+12\pi}{6}\] \[\Large -\frac{5\pi}{6} + 2\pi = \Large \frac{7\pi}{6}\]

jimthompson5910 (jim_thompson5910):

notice how I did not approximate pi and left it alone

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

both angles are in quadrant 3 

OpenStudy (anonymous):

O yes that makes sense. how would I do the opposite like if had to find something coterminal with 675 degrees?

jimthompson5910 (jim_thompson5910):

keep subtracting 360 degrees from your angle until you end up in the interval [0, 360)

jimthompson5910 (jim_thompson5910):

degrees are easier to work with in my opinion since you often don't have fractions and you won't have to deal with pi

OpenStudy (anonymous):

Yes degrees are much easier! Thank you for your help

jimthompson5910 (jim_thompson5910):

you're welcome